OpenStudy (anonymous):

Show that the set of complex numbers z which satisfy |z| = |z - 1 +2i|, (where i = sqrt(-1)), is a line in the argand plane and sketch this line. Please show steps. Many thanks

OpenStudy (helder_edwin):

let \(z=a+bi\) then \[ \large |z|=|z-1+2i| \] \[ \large |a+bi|=|a+bi-1+2i|=|(a-1)+(b+2)i| \] \[ \large \sqrt{a ^2+b^2}=\sqrt{(a-1)^2+(b+2)^2} \]

OpenStudy (helder_edwin):

do u get this?

OpenStudy (anonymous):

One second please

OpenStudy (anonymous):

Do you mind explainig please

OpenStudy (helder_edwin):

since u dont know z u make it ANY complex number that is the first sentence

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

now u substitute this into the equation and apply the definition of modulus of a complex number

OpenStudy (anonymous):

I understand, yes

OpenStudy (anonymous):

is there more to follow?

OpenStudy (helder_edwin):

can u go on by yourself?

OpenStudy (anonymous):

I can do the arithmetic, no problem but I have two questions 1)do i replace a and b with integers. 2) And do sketch a normal argand diagram?

OpenStudy (helder_edwin):

1) NO. a and b are just variables. u could have just as well used x and y. plus u are asked for an equation. 2) the argand plane is just the cartesian plane. remember \(\mathbb{C}=\mathbb{R}^2\)

OpenStudy (helder_edwin):

tell me whay u get

OpenStudy (anonymous):

one minute please

OpenStudy (anonymous):

I am a bit stuck\[\sqrt{a ^{2}+b ^{2}}=\sqrt{a ^{2}-2a+b{^2}+4b+4}\] is what I get

OpenStudy (helder_edwin):

OK let's do it first u cancel both square roots \[ \large a^2+b^2=(a-1)^2+(b+2)^2 \] now expand \[ \large a^2+b^2=a^2-2a+1+b^2+4b+4 \] now cancel like items \[ \large 0=-2a+1+4b+4=-2a+4b+5 \]

OpenStudy (helder_edwin):

understand?

OpenStudy (anonymous):

okay

OpenStudy (helder_edwin):

this is your equation

OpenStudy (anonymous):

that is the 1st part many thanks

OpenStudy (anonymous):

but how would you sketch that on the argand plane, since it does not have an imaginary part

OpenStudy (helder_edwin):

remember \(z=a+bi\)

OpenStudy (helder_edwin):

a is the real part and b is the imaginary part

OpenStudy (anonymous):

okay and the 5?

OpenStudy (helder_edwin):

it is just a constant u can rewrite the equation into the form \[ \large b=ma+c \]

OpenStudy (helder_edwin):

like in analytic geometry

OpenStudy (anonymous):

okay , nice one, helder_edwin, under which specific section would I find this topic should I wish to read further? You have been great - an officer and a gentlemen!

OpenStudy (helder_edwin):

every book on complex analysis has a first chapter about complex numbers

OpenStudy (helder_edwin):

thank you for your words

OpenStudy (anonymous):

okay, I have a 1st year (university) engineering maths book, I will go through it again to search for this particular problem. I will write you a good testimonial. Thank you so much once and goodbye.

OpenStudy (helder_edwin):

this might help http://uploading.com/files/5abe2cc8/3540939822Complex%2BAnalysisB.rar/

OpenStudy (anonymous):

DL is a bit tricky , its asking me to dl an app? Is that correct?

OpenStudy (helder_edwin):

on the left of the screen there is a counter stopped at 60 and a button that says start timer press it it will start counting down when it reaches zero a window will pop up to ask u what to do with the file

OpenStudy (helder_edwin):

or if u want, i can mail the book to u

OpenStudy (anonymous):

Thanks for the offer - Its busy DLing right now. It was more tricky than usual :) How long have you been on open study helder?

OpenStudy (helder_edwin):

i dont know less than a month how about u?

OpenStudy (anonymous):

I registered some months back but just started using it a lot about two weeks ago. Its really great and what fascinates me is that the help is very fast and free! Unbelievable !

OpenStudy (helder_edwin):

yeah it a great place

OpenStudy (anonymous):

Well helder, I am knackered and want to retire to bed. Its 3am here in Galway, West of Ireland, so I need to get some shut eye. Are you Stateside?

OpenStudy (helder_edwin):

i am in Bolivia, south america

OpenStudy (helder_edwin):

rest well it was nice meeting you

OpenStudy (anonymous):

Wow, lovely. Well I hope to catch you again sometime here on openstudy - compadre. Keep well. Adios Amigo. :)

OpenStudy (helder_edwin):

adios